OpenStudy (anonymous):

If a, b are the solutions of (((15)/(t+6))+((t+4)/(3)))=4 compute the sum of a+b plz show work and i would be very happy if someone were to help me/explain this problem

OpenStudy (anonymous):

first let me see if i can read what is says

OpenStudy (anonymous):

\[\frac{15}{t+6}+\frac{t+4}{3}=4\]??

OpenStudy (anonymous):

yuppers

OpenStudy (anonymous):

the sum of a+b for tht equation

OpenStudy (anonymous):

well i don't think there is a shortcut. i think you have to find the solutions and add them

OpenStudy (anonymous):

or at least write the quadratic equation out

OpenStudy (anonymous):

oh damn i made a mistake!

OpenStudy (anonymous):

\[\frac{45+(t+4)(t+6)}{3(t+6)}=4\]

OpenStudy (anonymous):

\[45+(t+4)(t+6)=12(t+6)\] \[t^2+10t+69=12t+72\] \[t^2-2t-3=0\] and finally \[(t-3)(t+1)=0\] solutions area 3 and -1

OpenStudy (anonymous):

if you add them you get 2

OpenStudy (anonymous):

it is also true that the sum of the solutions to \[ax^2+bx+c=0\] is \[-\frac{b}{a}\]so we did not actually have to solve once we got to \[x^2-2x-3=0\]

OpenStudy (anonymous):

thank much man appreciate it hope some good karma comes back to u one day ^^